OpenStudy (anonymous):

Find dy/dx given that y² =x(3-x)². State the values of x for which dy/dx is: (a) zero (b) infinite. answer to (a) ± (3/2√x-(3/2)√x) answer to (b) is 1 HI CAN YOU SHOW ME STEP BY STEP HOW TO GET X=± (3/2√x-(3/2)√x). AS AM GETTING A DIFFERENT ANSWER (x^2=9/12x) PS and part b how do i get the 1 thanks

OpenStudy (shamil98):

is that a z in the exponent or a 2?

OpenStudy (anonymous):

its 2

OpenStudy (shamil98):

Ok.

OpenStudy (shamil98):

Show your attempt at finding dy/dx first.

OpenStudy (anonymous):

ok using implicit differentiation

OpenStudy (anonymous):

and also function of a function rule for (3-x)²

OpenStudy (anonymous):

i get 2ydy/dx=-2x (3-x)²+ (3-x)².1

OpenStudy (anonymous):

sorry it should be 2ydy/dx= 2x (3-x)+(3-x)²

OpenStudy (anonymous):

when you expand and set dy/dx=0 you get 0=-12x-x²+9

OpenStudy (anonymous):

getting x²=9/12x

OpenStudy (anonymous):

which doesn't seem correct

OpenStudy (shamil98):

y² =x(3-x)² Using the chain rule and the product rule together: \[ 2yy' = 1*x(3-x)^2 + 2x(3-x) * (-1) \] \[\ y' = \frac {x(3-x)^2 - 2x(3-x)} {2} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x(3-x)^2 (x)(2(3-x)*-1)+(1(3-x)^2) (x)(6-2x-1)+(3-x)^2 6x-2x^2-x +(3-x)^2 -2x^2+5x+(3-x)^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

expanding (3-x)^2 we get 9-6x+x^2

OpenStudy (anonymous):

so we get -x^2-x+9

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

Using the product rule and chain rule, yes. I am trying to figure out why shamil has x's in front of the parenthesis

OpenStudy (neer2890):

I think shamil has written x by mistaken

OpenStudy (neer2890):

also, in 2nd step, denominator should be 2y

OpenStudy (anonymous):

\[y'=\frac{ 3(x^2-4x+3) }{ 2y }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so do we substitute y into the denominator?